Identify the focus and directrix of the following parabola
(a) 2y^2 = x
(b) x^2 = -36y
1
Expert's answer
2021-01-28T23:59:12-0500

a) "4p(x-h)=(y-k)^2"

Focus is "(h+p,k)"

Directrix is "x=h-p"

We have:

"p=1\/8, h=k=0"

focus: "(1\/8,0)"

directrix: "x=-1\/8"


b) "4p(y-k)=(x-h)^2"

Focus is "(h,k+p)"

Directrix is "y=k-p"

We have:

"p=-9, h=k=0"

focus: "(0,-9)"

directrix: "y=9"